# Arm: Adaptive Reasoning Model

### v1

@@ 1 Introduction
The emergence of large reasoning models (LRMs) such as OpenAI-o1 [16] and DeepSeek-R1 [9] has led to unprecedented breakthroughs in problem-solving capabilities through test-time scaling [3; 51]. These models are designed to solve tasks using Long Chain-of-Thought (Long CoT), generating more tokens to achieve better performance. However, because they are primarily trained on tasks requiring intensive reasoning, LRMs tend to apply Long CoT uniformly across all tasks, resulting in the so-called “overthinking” problem [4; 37]. This issue refers to the excessive use of tokens for reasoning, which yields no performance gains and may introduce noise that misleads the model [46; 7].
While some efforts aim to reduce token usage in LRMs, they often rely on clear estimations of the token budget per task [1; 41] or require specialized, length-constrained model training [13]. In reality, such estimations are not always accurate, and a more desirable solution is for models to adaptively control their token usage based on task complexity without human intervention. For example, for the first easy question in Figure 1(a), answering directly is the ideal choice, whereas for the second, hard question, the use of Long CoT is precisely what is needed. Thus, Long CoT is not a “silver bullet”; selecting an appropriate reasoning format is essential to balance efficiency and effectiveness.
@@ 2.1 Reinforcement Learning for Improving Reasoning
Reinforcement Learning (RL) has demonstrated significant potential in enhancing the problem-solving abilities of large language models (LLMs) across various domains [29; 42; 17]. Recently, Reinforcement Learning with Verifiable Rewards (RLVR) has gained substantial attention for advancing LLM capabilities [18; 9; 21], resulting in the development of large reasoning models (LRMs) [47] such as OpenAI-o1 [16] and DeepSeek-R1 [9]. Based on simple rule-based rewards, RLVR algorithms such as Group Relative Policy Optimization (GRPO) [33] enable models to use Long Chain-of-Thought (Long CoT) [50; 15]. This facilitates deep reasoning behaviors, such as searching, backtracking, and verifying through test-time scaling [3; 51]. However, these models also suffer from significant computational overhead due to extended outputs across all tasks, leading to inefficiency associated with the “overthinking” phenomenon [4; 31; 37]. Verbose and redundant outputs can obscure logical clarity and hinder the model’s ability to solve problems effectively [46; 7].
@@ 2.2 Efficiency in Large Language Models
Recently, many studies have focused on improving the reasoning efficiency in LLMs. Some prompt-guided methods [10; 48; 19] explicitly instruct LLMs to generate concise reasoning outputs by controlling input properties such as task difficulty and response length. Other approaches [11; 5; 34] explore training LLMs to reason in latent space, generating the direct answer without the need for detailed language tokens. Several techniques have also been proposed to reduce inference costs by controlling or pruning output length, either by injecting multiple reasoning formats during the pre-training stage [36] or by applying length penalties during the RL stage [40; 2; 1; 13]. Many of these methods aim to strike a trade-off between token budget and reasoning performance by shortening output lengths, often relying on clear estimations of the token budget for each task or requiring specialized, length-constrained model training. However, in reality, such estimations are not always accurate, and what we ultimately expect is for models to adaptively regulate their token usage based on the complexity of the task at hand. Therefore, in this work, we propose a novel training framework that enables models to adaptively select suitable reasoning formats for given tasks by themselves, optimizing both performance and computational efficiency.
@@ 3.1 Stage 1: SFT for Reasoning Formats Understanding
In this stage, we leverage SFT as a cold start to introduce the model to various reasoning formats it can utilize to solve problems.111In preliminary experiments, models without SFT failed to distinguish between the four reasoning formats, like producing mixed outputs, wrapping a Short CoT response using the special tokens intended for Long CoT. These formats include three efficient reasoning formats Direct Answer, Short CoT, and Code, as well as the elaborate reasoning format Long CoT. We use special tokens (e.g., <Code></Code>) to embrace thinking rationale. Specifically, 1) Direct Answer:This format provides a direct answer without any reasoning chain, making it the most efficient in terms of token usage. 2) Short CoT:This format begins with a short reasoning and then provides an answer, which has been proved effective in mathematical problems [43]. 3) Code:This format adopts code-based reasoning, which has proven effective across a variety of tasks due to its structured process [44; 45; 20]. 4) Long CoT:This format involves a more detailed, iterative reasoning process, thus incurs higher token usage. It is suited for tasks requiring advanced reasoning capabilities, such as self-reflection and alternative generation, where those more efficient formats fall short [27; 9; 50].
@@ 4.1 Experimental Setup
Model To assess the effectiveness of our method across models of different sizes, we select Qwen2.5-Base-3B/7B/14B [49] as backbone models. We further examine models of the same family but fine-tuned on different datasets, specifically the Instruct [49] and DeepSeek-R1-Distill variants [9], which exhibit varying levels of base reasoning capabilities. A detailed analysis is given in Section 5.3.
Training Datasets Stage 1: We use AQuA-Rat [22] as the SFT dataset, as its answers can be naturally transformed into four distinct reasoning formats. In addition to the Direct Answer and Short CoT rationales provided with the dataset, we utilize GPT-4o [26] and DeepSeek-R1 [9] to supplement the Code and Long CoT rationales, respectively. To ensure the quality of the generated rationales, we filter out those that lead to incorrect answers, resulting in a training set containing K multiple-choice and K open-form questions, each with four reasoning formats. Appendix B provides further details on the generation and filtering process. Stage 2: To prevent data leakage, we employ three additional datasets exclusively for the RL stage.222In preliminary experiments, we observed that using the same training data in both stages causes the model to recite answers rather than reasoning during the RL stage, resulting in poor generalization. These datasets cover a range of difficulty levels, from relatively simple commonsense reasoning tasks to more complex mathematical reasoning tasks, including CommonsenseQA (CSQA) [39], GSM8K [6], and MATH [12], collectively comprising K verifiable question-answer pairs. Please refer to Appendix C for details of the datasets we used and Appendix D for implementation details.
Baselines In addition to backbone models, we compare Arm with models trained using alternative algorithms that may enable adaptive reasoning capabilities. Specifically, Qwen2.5 refers to the backbone model trained on the AQuA-Rat dataset used in stage 1. In this setting, we explore whether language models can master adaptive reasoning through a straightforward SFT strategy. For Qwen2.5, we examine whether SFT models, further trained with GRPO, can better understand different reasoning formats and whether this approach empowers them to select appropriate reasoning formats based on rule-based rewards.
@@ 4.3 Main Results
GRPO does improve reasoning capabilities, but it tends to rely on Long CoT to solve all tasks. We observe that models trained with GRPO achieve significant improvements across all tasks, yet the token cost remains substantial, especially for the two easier tasks. Further analysis reveals that Long CoT is predominantly used in the inference stage, as shown in Figure 2. This behavior stems from the nature of GRPO (i.e., format collapse discussed in Section 3.2), where models converge to the format with the highest accuracy (i.e., Long CoT) early in training (10 steps in our experiment). As a result, GRPO also fails to teach models how to select a more efficient reasoning format based on the task.
Arm is able to adaptively select reasoning formats based on task difficulty, while achieving comparable accuracy across all tasks compared to GRPO and using significantly fewer tokens. As shown in Table 1, across three different model sizes, all Arms experience an average performance drop of less than 1% compared to models trained with GRPO, yet they save more than 30% of the tokens. Specifically, Arm demonstrates a clear advantage on easy tasks, saving over 70% of tokens while maintaining comparable accuracy. This advantage extends to medium tasks as well. For the more challenging AIME’25 task, Arm adapts to the task difficulty by increasingly selecting Long CoT, thereby avoiding performance degradation on harder tasks, with Arm-14B even surpassing its counterpart Qwen2.5-14B. Figure 2 further confirms that Arm is able to gradually adopt more advanced reasoning formats and discards simpler ones as task difficulty increases. Moreover, as shown in Figure 1(b), the line connecting “SFT” and “+GRPO” illustrates the expected trade-off, while “+Ada-GRPO” consistently lies above it, indicating a better balance between effectiveness and efficiency of Arm. Additionally, Arm-7B achieves comparable performance to DS-R1-Distill-7B while using only % of the tokens on average.
@@ 4.4 Reasoning Mode Switching
To evaluate the performance and effectiveness of the proposed reasoning modes, we conduct experiments across various evaluation datasets. Table 2 presents the results for Arm-7B. Specifically: 1) Adaptive Mode strikes a superior balance between high accuracy and efficient token usage across all datasets, demonstrating its ability to adaptively select the reasoning formats. 2) Instruction-Guided Mode offers a clear advantage when the assigned reasoning format is appropriate.For example, Direct Answer is sufficient for commonsense tasks, while Code, due to its structured nature, performs better on symbolic reasoning tasks compared to Direct Answer and Short CoT. Furthermore, Inst achieves better performance (74.5%) than the same-sized model trained on GRPO (73.2% in Table 1). This demonstrates that Ada-GRPO does not hinder the model’s Long CoT reasoning capabilities. We further validate this by analyzing the reflective words used by Arm-7B and Qwen2.5-7B in Appendix E. 3) Consensus-Guided Mode, on the other hand, is performance-oriented, requiring more tokens to achieve better performance.This mode leverages consensus across multiple formats to mitigate bias and uncertainty present in any single format, offering greater reliability, particularly for reasoning tasks that demand advanced cognitive capabilities, where simpler formats may fall short. This is evidenced by the fact that Long CoT is less likely to be used for easy tasks, but is highly likely to be selected for medium tasks and even used 100% of the time for the most difficult AIME’25 task.
@@ 5.1 Effectiveness of Adaptive Format Selection
To verify that Arm’s format selection indeed adapts to the task at hand rather than relying on random selection, we compare Arm’s Adaptive Mode with Instruction-Guided Mode. In Instruction-Guided Mode, the reasoning format is fixed and manually specified, providing a strong baseline to test whether adaptive selection offers real benefits over using a uniform format across tasks. We report the accuracy of both modes in Figure 3. We observe that the accuracy of the reasoning formats selected in Adaptive Mode is higher than that in Instruction-Guided Mode. Specifically, Adaptive Mode improves accuracy by 4.7% on Direct Answer, by 2.7% on both Short CoT and Code, and even yields a slight improvement on Long CoT. These results confirm that Arm is not randomly switching formats but is instead learning to select an appropriate one for each task.
@@ 5.3 Comparison of Backbone Models
Beyond the base model, we further analyze the impact of different backbone models, including instruction-tuned and DS-R1-Distill variants. Figure 6 reports accuracy and token usage across easy, medium, and hard tasks. We observe that base and instruction-tuned models have a highly similar performance. This suggests that RL effectively bridges the gap left by instruction tuning, enabling base models to achieve comparable performance, consistent with findings from previous work [17]. In contrast, the DS-R1-Distill variant performs notably better on medium and hard tasks, benefiting from distilled knowledge from the stronger DeepSeek-R1 model, though at the expense of increased token cost. However, it performs significantly worse on easy tasks, even with excessive token usage, resulting from the overthinking phenomenon. For more discussion and case studies on the overthinking phenomenon, please refer to Appendix F.
@@ 5.4 Comparison of Arm and Length-Penalty-Based Strategies
To examine whether previously proposed length-penalty-based strategies—proven effective in complex reasoning—remain effective for easier tasks, we evaluate two representative methods, L1 [1] and ThinkPrune[13], on the CSQA dataset. Since both methods are based on the DS-R1-Distill model, we ensure a fair comparison by also evaluating the version of Arm trained on the same backbone. We report the relative accuracy and token usage of all three models compared to their respective backbone models in Figure 4. When using the minimum allowed lengths specified in the official settings of L1 and ThinkPrune, both methods exhibit performance drops. In contrast, Arm maintains strong performance while using relatively fewer tokens, demonstrating its ability to balance reasoning efficiency and effectiveness. Please refer to Appendix G for more details.
@@ D.2 Stage 2: RL
Appendix E Details of Reflective Words

### v2

@@ 1 Introduction
The emergence of large reasoning models (LRMs) such as OpenAI-o1 [19] and DeepSeek-R1 [11] has led to unprecedented breakthroughs in problem-solving capabilities through test-time scaling [3; 57]. These models are designed to solve tasks using Long Chain-of-Thought (Long CoT), generating more tokens to achieve better performance. However, because they are primarily trained on tasks requiring intensive reasoning, LRMs tend to apply Long CoT uniformly across all tasks, resulting in the so-called “overthinking” problem [4; 42]. This issue refers to the excessive use of tokens for reasoning, which yields no performance gains and may introduce noise that misleads the model [52; 7].
While some efforts aim to reduce token usage in LRMs, they often rely on clear estimations of the token budget per task [1; 46] or require specialized, length-constrained model training [16]. In reality, such estimations are not always accurate, and a more desirable solution is for models to adaptively control their token usage based on task complexity without human intervention. For example, for the first easy question in Figure 1(a), answering directly is the ideal choice, whereas for the second, hard question, the use of Long CoT is precisely what is needed. Thus, Long CoT is not a “silver bullet”; selecting an appropriate reasoning format is essential to balance efficiency and effectiveness.
@@ 2.1 Reinforcement Learning for Improving Reasoning
Reinforcement Learning (RL) has demonstrated significant potential in enhancing the problem-solving abilities of large language models (LLMs) across various domains [33; 47; 20]. Recently, Reinforcement Learning with Verifiable Rewards (RLVR) has gained substantial attention for advancing LLM capabilities [21; 11; 25], resulting in the development of large reasoning models (LRMs) [53] such as OpenAI-o1 [19] and DeepSeek-R1 [11]. Based on simple rule-based rewards, RLVR algorithms such as Group Relative Policy Optimization (GRPO) [38] enable models to use Long Chain-of-Thought (Long CoT) [56; 18]. This facilitates deep reasoning behaviors, such as searching, backtracking, and verifying through test-time scaling [3; 57]. However, these models also suffer from significant computational overhead due to extended outputs across all tasks, leading to inefficiency associated with the “overthinking” phenomenon [4; 35; 42]. Verbose and redundant outputs can obscure logical clarity and hinder the model’s ability to solve problems effectively [52; 7].
@@ 2.2 Efficiency in Large Language Models
Recently, many studies have focused on improving reasoning efficiency in LLMs. Some prompt-guided methods [12; 13; 54; 22] explicitly instruct LLMs to generate concise reasoning outputs by controlling input properties such as task difficulty and response length. Other approaches [14; 5; 39] explore training LLMs to reason in latent space, generating the direct answer without the need for detailed language tokens. Several techniques have also been proposed to reduce inference costs by controlling or pruning output length, either by injecting multiple reasoning formats during the pre-training stage [41] or by applying length penalties during the RL stage [45; 2; 1; 16]. Many of these methods aim to strike a trade-off between token budget and reasoning performance by shortening output lengths, often relying on clear estimations of the token budget for each task or requiring specialized, length-constrained model training. However, in reality, such estimations are not always accurate, and what we ultimately expect is for models to adaptively regulate their token usage based on the complexity of the task at hand. Therefore, in this work, we propose a novel training framework that enables models to adaptively select suitable reasoning formats for given tasks by themselves, optimizing both performance and computational efficiency.
@@ 3.1 Stage 1: SFT for Reasoning Formats Understanding
In this stage, we leverage SFT as a cold start to introduce the model to various reasoning formats it can utilize to solve problems.111In preliminary experiments, models without SFT failed to distinguish between the four reasoning formats, like producing mixed outputs, wrapping a Short CoT response using the special tokens intended for Long CoT. These formats include three efficient reasoning formats Direct Answer, Short CoT, and Code, as well as the elaborate reasoning format Long CoT. We use special tokens (e.g., <Code></Code>) to embrace thinking rationale. Specifically, 1) Direct Answer:This format provides a direct answer without any reasoning chain, making it the most efficient in terms of token usage. 2) Short CoT:This format begins with a short reasoning and then provides an answer, which has been proved effective in mathematical problems [49]. 3) Code:This format adopts code-based reasoning, which has proven effective across a variety of tasks due to its structured process [50; 51; 24]. 4) Long CoT:This format involves a more detailed, iterative reasoning process, thus incurs higher token usage. It is suited for tasks requiring advanced reasoning capabilities, such as self-reflection and alternative generation, where those more efficient formats fall short [31; 11; 56].
@@ 4.1 Experimental Setup
Model To assess the effectiveness of our method across different model sizes, we select Qwen2.5-Base-3B/7B/14B [55] as backbone models. We further examine models of the same family but fine-tuned on different datasets, specifically the Instruct [55] and DeepSeek-R1-Distill variants [11], which exhibit varying levels of base reasoning capabilities. A detailed analysis is given in Section 5.3.
Training Datasets Stage 1: We use AQuA-Rat [26] as the SFT dataset, as its answers can be naturally transformed into four distinct reasoning formats. In addition to the Direct Answer and Short CoT rationales provided with the dataset, we utilize GPT-4o [30] and DeepSeek-R1 [11] to supplement the Code and Long CoT rationales, respectively. To ensure the quality of the generated rationales, we filter out those that lead to incorrect answers, resulting in a training set containing K multiple-choice and K open-form questions, each with four reasoning formats. Appendix B provides further details on the generation and filtering process. Stage 2: To prevent data leakage, we employ three additional datasets exclusively for the RL stage.222In preliminary experiments, we observed that using the same training data in both stages causes the model to recite answers rather than reasoning during the RL stage, resulting in poor generalization. These datasets cover a range of difficulty levels, from relatively simple commonsense reasoning tasks to more complex mathematical reasoning tasks, including CommonsenseQA (CSQA) [44], GSM8K [6], and MATH [15], collectively comprising K verifiable question-answer pairs. Please refer to Appendix C for details of the datasets we use and Appendix D for implementation details.
Baselines In addition to backbone models, we compare Arm with models trained using alternative algorithms that may enable adaptive reasoning capabilities. Specifically, Qwen2.5 refers to the backbone models trained on the AQuA-Rat dataset used in Stage 1. In this setting, we explore whether language models can master adaptive reasoning through a straightforward SFT strategy. For Qwen2.5, we examine whether SFT models, further trained with GRPO, can better understand different reasoning formats and whether this approach empowers them to select appropriate reasoning formats based on rule-based rewards.
@@ 4.3 Main Results
GRPO does improve reasoning capabilities, but it tends to rely on Long CoT to solve all tasks. We observe that models trained with GRPO achieve significant improvements across all tasks, yet the token cost remains substantial, especially for the two easier tasks. Further analysis reveals that Long CoT is predominantly used in the inference stage, as shown in Figure 2. This behavior stems from the nature of GRPO (i.e., format collapse discussed in Section 3.2), where models converge to the format with the highest accuracy (i.e., Long CoT) early in training (10 steps in our experiment). As a result, GRPO also fails to teach models how to select a more efficient reasoning format based on the task. We provide more details of format collapse in Appendix E.
Arm is able to adaptively select reasoning formats based on task difficulty, while achieving comparable accuracy across all tasks compared to GRPO and using significantly fewer tokens. As shown in Table 1, across three different model sizes, all Arms experience an average performance drop of less than 1% compared to models trained with GRPO, yet they save more than 30% of the tokens. Specifically, Arm demonstrates a clear advantage on easy tasks, saving over 70% of tokens while maintaining comparable accuracy. This advantage extends to medium tasks as well. For the more challenging AIME’25 task, Arm adapts to the task difficulty by increasingly selecting Long CoT, thereby avoiding performance degradation on harder tasks, with Arm-14B even surpassing its counterpart Qwen2.5-14B. Figure 2 further confirms that Arm is able to gradually adopt more advanced reasoning formats and discards simpler ones as task difficulty increases. Moreover, as shown in Figure 1(b), the line connecting “SFT” and “+GRPO” illustrates the expected trade-off, while “+Ada-GRPO” consistently lies above it, indicating a better balance between effectiveness and efficiency of Arm. Additionally, Arm-7B achieves comparable performance to DS-R1-Distill-7B while using only % of the tokens on average. For a broader view of generalization across additional benchmarks, please refer to Appendix F.
@@ 4.4 Reasoning Mode Switching
To evaluate the performance and effectiveness of the proposed reasoning modes, we conduct experiments across various evaluation datasets. Table 2 presents the results for Arm-7B. Specifically: 1) Adaptive Mode strikes a superior balance between high accuracy and efficient token usage across all datasets, demonstrating its ability to adaptively select the reasoning formats. 2) Instruction-Guided Mode offers a clear advantage when the assigned reasoning format is appropriate.For example, Direct Answer is sufficient for commonsense tasks, while Code, due to its structured nature, performs better on symbolic reasoning tasks compared to Direct Answer and Short CoT. Furthermore, Inst achieves better performance (74.5%) than the same-sized model trained on GRPO (73.2% in Table 1). This demonstrates that Ada-GRPO does not hinder the model’s Long CoT reasoning capabilities. We further validate this by analyzing the reflective words used by Arm-7B and Qwen2.5-7B in Appendix G. 3) Consensus-Guided Mode, on the other hand, is performance-oriented, requiring more tokens to achieve better performance.This mode leverages consensus across multiple formats to mitigate bias and uncertainty present in any single format, offering greater reliability, particularly for reasoning tasks that demand advanced cognitive capabilities, where simpler formats may fall short. This is evidenced by the fact that Long CoT is less likely to be used for easy tasks, but is highly likely to be selected for medium tasks and even used 100% of the time for the most difficult AIME’25 task.
@@ 5.1 Effectiveness of Adaptive Format Selection
To verify that Arm’s format selection indeed adapts to the task at hand rather than relying on random selection, we compare Arm’s Adaptive Mode with Instruction-Guided Mode. In Instruction-Guided Mode, the reasoning format is fixed and manually specified, providing a strong baseline to test whether adaptive selection offers real benefits over using a uniform format across tasks. We report the accuracy of both modes in Figure 3. We observe that the accuracy of the reasoning formats selected in Adaptive Mode is higher than that in Instruction-Guided Mode. Specifically, Adaptive Mode improves accuracy by 4.7% on Direct Answer, by 2.7% on both Short CoT and Code, and even yields a slight improvement on Long CoT. These results confirm that Arm is not randomly switching formats but is instead learning to select an appropriate one for each task. Further ablation results on reasoning formats are presented in Appendix H.
@@ 5.3 Comparison of Backbone Models
Beyond the base model, we further analyze the impact of different backbone models, including instruction-tuned and DS-R1-Distill variants. Figure 6 reports accuracy and token usage across easy, medium, and hard tasks. We observe that base and instruction-tuned models have a highly similar performance. This suggests that RL effectively bridges the gap left by instruction tuning, enabling base models to achieve comparable performance, consistent with findings from previous work [20]. In contrast, the DS-R1-Distill variant performs notably better on medium and hard tasks, benefiting from distilled knowledge from the stronger DeepSeek-R1 model, though at the expense of increased token cost. However, it performs significantly worse on easy tasks, even with excessive token usage, resulting from the overthinking phenomenon. Additional discussion and case studies on the overthinking phenomenon are presented in Appendix I, and a complementary analysis of LLaMA-based backbones is included in Appendix J.
@@ 5.4 Comparison of Arm and Length-Penalty-Based Strategies
To examine whether previously proposed length-penalty-based strategies—proven effective in complex reasoning—remain effective for easier tasks, we evaluate two representative methods, L1 [1] and ThinkPrune [16], on the CSQA dataset. Since both methods are based on the DS-R1-Distill model, we ensure a fair comparison by also evaluating the version of Arm trained on the same backbone. We report the relative accuracy and token usage of all three models compared to their respective backbone models in Figure 4. When using the minimum allowed lengths specified in the official settings of L1 and ThinkPrune, both methods exhibit performance drops. In contrast, Arm maintains strong performance while using relatively fewer tokens, demonstrating its ability to balance reasoning efficiency and effectiveness. Please see Appendix K for further discussion and details.
@@ D.2 Stage 2: RL
Appendix E Effect of Training Data Bias on Format Collapse
Since training data may implicitly favor certain reasoning formats, it is important to examine whether such bias can induce format collapse. To this end, we analyze an Arm trained solely on the AIME dataset (1983-2024), which primarily favors Long CoT solutions due to its competition-level complexity. We present the result in Table 4. We evaluate this model on two simpler tasks—CSQA and OBQA—and observe that the model overwhelmingly selects Long CoT (80%) even when simpler formats would suffice. This confirms that training on a biased dataset can indeed lead to over-reliance on a single reasoning format.
| 7B Models | CSQA | OBQA | ||
| Long CoT | Other Formats | Long CoT | Other Formats | |
| Training with AIME only | 79.4% | 20.6% | 83.0% | 17.0% |
To mitigate this, Arm is trained on a diverse mixture of datasets across a wide range of difficulties. As shown in Table 5, the full Arm recipe achieves both higher accuracy and significantly reduced token usage on the tasks, demonstrating that our approach effectively prevents format collapse and encourages adaptive reasoning behavior across domains.
| 7B Models | CSQA | OBQA | ||
| Acc. | Tok. | Acc. | Tok. | |
| Training with AIME only | 78.6 | 401 | 82.2 | 426 |
| Arm Recipe | 86.1 | 136 | 84.4 | 159 |
Appendix F Generalization to Additional Benchmarks
As shown in Section 4.2, our existing evaluation spans across in- and out-of-domain commonsense, mathematical, and symbolic reasoning. To further investigate the generalizability of Arm, we extend the evaluation to two additional benchmarks: 1) GPQA[Rein et al., 2024], a challenging QA dataset designed to test compositional reasoning, and 2) StrategyQA[Geva et al., 2021], a benchmark of yes/no questions that require implicit multi-hop reasoning.
As shown in Table 6, Arm achieves comparable accuracy to the GRPO baseline while significantly reducing token cost by an average of 50%, and up to 65% in StrategyQA, consistent with our main findings discussed in Section 4.3. This demonstrates that Arm’s adaptive reasoning behavior generalizes well to more tasks.
| Models | GPQA-Main | GPQA-Diamond | StrategyQA | |||
| Acc. | Tok. | Acc. | Tok. | Acc. | Tok. | |
| Qwen2.5-7B | 35.0 | 2324 | 37.4 | 2604 | 72.9 | 646 |
| Arm-7B | 34.8 | 1306 | 36.9 | 1536 | 73.8 | 229 |
Appendix G Details of Reflective Words